The section of 2x^2+y^2=2(1-z^2) by the plane x+2=0 is a hyerbola.

True or false with full explanation
1
Expert's answer
2020-12-03T07:21:41-0500

The general equation of a hyperbola is


"\frac{x^2}{a^2}-\frac{y^2}{b^2}=1".


Let us find the section of "2x^2+y^2=2(1-z^2)" by the plane "x+2=0":


"2(-2)^2+y^2=2(1-z^2)"


"8+y^2=2-2z^2"


"y^2+2z^2=-6"


Since the eqution has no real solution, the plane "x+2=0" does not intersect "2x^2+y^2=2(1-z^2)". The equation "y^2+2z^2=-6" is not a hyperbola equation.


Answer: false